OpenStudy (anonymous):

what would this look like? The difference of a number squared and the number is one less than four times the number.

hero (hero):

\(n^2 - n = 4n - 1\)

OpenStudy (anonymous):

thank you @hero

hero (hero):

You still have to solve for n right?

OpenStudy (anonymous):

"difference of" : Think minus "a number squared" : n^2 "is" : Think = " one less than 4 times the number": 4n - 1

OpenStudy (anonymous):

yes i do

hero (hero):

Are you familiar with the quadratic formula?

OpenStudy (anonymous):

yes I got 1/2? am I right?

hero (hero):

There are two values for n, right?

hero (hero):

But 1/2 isn't one of them

OpenStudy (anonymous):

um I dont know then?

hero (hero):

You've never seen this before \[n = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

no

hero (hero):

Are you sure you're required to solve for \(n\)? You won't be able to solve for \(n\) unless you are familiar with that.

hero (hero):

It is derived from the general form of a quadratic equation \(y = ax^2 + bx + c\)

OpenStudy (anonymous):

maybe im not supposed to all my teacher gave me was that wording?